Solve the following pair of equations graphically
2x - 3y = 1
1.25x + 0.5y = 3

Answer by Theo(13342)

the graph is shown below.
the intersection point is (x,y = (2,1).



these are two equations that need to be solved simultaneously.
one way to do that is shown below.

start with:

2x - 3y = 1
1.25x + .5y = 3

multiply both sides of the second equation by 6 and leave the first equation as is to get:

2x - 3y = 1
7.5x + 3y = 18

add the two equations together to get:
9.5x = 19
solve for x to get:
x = 19/9.5 = 2

in the first equation, replace x with 2 to get:
2x - 3y = 1 becomes 4 - 3y = 1.
add 3y to both sides of the equation and subtract 1 from both sides of the equation to get:
4 - 1 = 3y
combine like terms to get:
3 = 3y
solve for y to get:
y = 3/3 = 1

your solution is that (x,y) = (2,1).
same solution as on the graph.